# Performance Analysis of Physical Layer Security: From Far-Field to Near-Field

### v1

@@ I Introduction
Recent research trends have highlighted the advantages of multiple-antenna technologies in improving secrecy capacity through beamforming design [1, 3]. Intuitively, using a larger-aperture antenna array can provide a higher beamforming gain and, consequently, a higher secure transmission rate. With advancements in antenna manufacturing and to meet the stringent security demands of future networks, there is a growing trend for base stations (BSs) to employ extremely large-aperture antenna arrays and operate at significantly higher frequencies [4, 5]. However, this shift represents not just a quantitative increase in antenna size and carrier frequency but also a qualitative paradigm shift from conventional far-field communications (FFC) to near-field communications (NFC). Specifically, based on the criterion for differentiating between near-field and far-field regions—the Rayleigh distance $\frac{2D^{2}}{\lambda}$, where $D$ represents the antenna aperture and $\lambda$ is the wavelength [6]—the near-field region is expected to expand substantially. For instance, an array with an aperture of $D=0.5$ m operating at a frequency of 60 GHz would encompass a near-field region extending up to $100$ meters.
@@
In a nutshell, the UPW model, commonly utilized in the far-field region, is inapplicable for NFC. Therefore, it becomes imperative to reevaluate secrecy performance from a near-field perspective, considering the implications of these propagation differences on secure transmission.
@@ I-A Prior Works
The secrecy performance for FFC, where the UPW model is applied, has been widely studied; see the review papers [1, 3, 8], and related references. However, research on near-field physical layer security is still in its early stages. The authors in [9] utilized the USW model to study a near-field secure directional modulation system and propose a secure precoding algorithm. The work in [10] leveraged the NUSW model to develop a two-stage algorithm that maximizes the near-field secrecy capacity via jointly optimizing unit-modulus phase shifters and baseband digital beamformers. The authors of [11] introduced analog beamfocusing techniques using true-time delayers and phase shifters to effectively mitigate beamsplit and exploit near-field advantages, significantly improving physical layer security in near-field wideband communications. Meanwhile, the study in [12] proposed an efficient low-complexity strategy for designing the beamforming with artificial noise to ensure near-field secure transmission with multiple legitimate users.
@@
I-B Contributions
Motivated by the above research gaps, this article analyzes the performance of physical layer security across commonly used far-field and near-field channel models, namely the UPW, USW, and NUSW models. We derive closed-form expressions for both the secrecy capacity and the minimum required power, facilitating a straightforward comparison between each model. Besides, asymptotic analysis is performed to provide deeper insights. Our main contributions are summarized as follows.
  • We investigate secure transmission scenarios where a large‐aperture array communicates with a legitimate user while being eavesdropped by a legitimate user. We provide detailed analysis of various channel models applicable across different ranges in both near-field and far-field regions, including the UPW, USW and NUSW models. In this context, we introduce two key metrics to measure secrecy performance: the secrecy capacity under a power constraint and the minimum power requirement needed to achieve a target secrecy rate.
@@ II System Model
We consider a multiple-input single-output (MISO) wiretap channel, which comprises a base station (BS), a legitimate user (Bob), and an eavesdropper (Eve), as depicted in Figure 2. The BS is equipped with a large-aperture UPA containing $M\gg 1$ antenna elements, while both users are single-antenna devices. The UPA is placed on the $x$-$z$ plane and centered at the origin. Here, $M=M_{x}M_{z}$, where $M_{x}$ and $M_{z}$ denote the number of array elements along the $x$- and $z$-axes, respectively. For brevity, we assume that $M_{x}$ and $M_{z}$ are odd numbers with $M_{x}=2\tilde{M}_{x}+1$ and $M_{z}=2\tilde{M}_{z}+1$. The physical dimensions of each BS array element along the $x$- and $z$-axes are denoted by $\sqrt{A}$, and the inter-element distance is $d$, where $d\geq\sqrt{A}$.
@@ II-A1 Secrecy Channel Capacity
${\mathsf{C}}=\max_{\lVert{\mathbf{w}}\rVert^{2}\leq P}{\mathsf{C}}_{\mathsf{w}}=\max_{\lVert{\mathbf{w}}\rVert^{2}\leq P}\left(\max\left\{\log_{2}\left(\frac{1+\gamma_{\mathsf{b}}}{1+\gamma_{\mathsf{e}}}\right),0\right\}\right).$ (4)
@@ II-B Channel Model
All channel gains are fixed throughout the entire transmission period, and each node $k\in\{{\mathsf{b}},{\mathsf{e}}\}$ is assumed to possess complete channel state information (CSI) about its effective channel. It is further assumed that Eve is a registered user, and hence, during the channel training phase, the BS acquires the CSI of both Eve and Bob. Under these circumstances, Eve is expected to receive the common messages broadcasted in the network but should remain uninformed about the confidential messages intended for Bob; see for example [17] and references therein.
As a theoretical exploration of fundamental performance limits and asymptotic behaviors, we consider basic free-space line-of-sight (LoS) propagation scenarios. Following our description concerning Figure 2, the central location of the $(m_{x},m_{z})$th BS element is denoted by ${\mathbf{s}}_{m_{x},m_{z}}=[m_{x}d,0,m_{z}d]^{\mathsf{T}}$, where $m_{x}\in{\mathcal{M}}_{x}\triangleq\{0,\pm 1,\ldots,\pm\tilde{M}_{x}\}$ and $m_{z}\in{\mathcal{M}}_{z}\triangleq\{0,\pm 1,\ldots,\pm\tilde{M}_{z}\}$. Consequently, the physical dimensions of the UPA along the $x$- and $z$-axes are $L_{x}\approx M_{x}d$ and $L_{z}\approx M_{z}d$, respectively.
Let $r_{k}$ denote the distance from the center of the antenna array to each node $k\in\{{\mathsf{b}},{\mathsf{e}}\}$, and $\theta_{k}\in(0,\pi)$ and $\phi_{k}\in(0,\pi)$ denote the azimuth and elevation angles, respectively. Thus, the location of node $k$ can be expressed as ${\mathbf{r}}_{k}=[r_{k}\Phi_{k},r_{k}\Psi_{k},r_{k}\Omega_{k}]^{\mathsf{T}}$, where $\Phi_{k}\triangleq\sin{\phi_{k}}\cos{\theta_{k}}$, $\Psi_{k}\triangleq\sin\phi_{k}\sin\theta_{k}$, and $\Omega_{k}\triangleq\cos{\phi_{k}}$. Moreover, we assume that both Bob and Eve are equipped with a single hypothetical isotropic antenna to receive incoming signals.
@@ II-B1 NUSW Model
$r_{m_{x},m_{z},k}=\lVert{\mathbf{r}}_{k}-{\mathbf{s}}_{m_{x},m_{z}}\rVert$ (6a)
$=r_{k}\sqrt{1-2m_{x}\varepsilon_{k}\Phi_{k}-2m_{z}\varepsilon_{k}\Omega_{k}+(m_{x}^{2}+m_{z}^{2})\varepsilon_{k}^{2}},$ (6b)
@@
$g_{m_{x},m_{z},k}=\int_{{\mathcal{S}}_{m_{x},m_{z}}}{\mathsf{g}}(\mathbf{r}_{k},{\mathbf{s}}){\frac{\lvert{\mathbf{e}}_{y}^{\mathsf{T}}({\mathbf{s}}-{\mathbf{r}}_{k})\rvert}{\lVert{\mathbf{r}}_{k}-{\mathbf{s}}\rVert}}{\rm{d}}{\mathbf{s}}.$ (7)
where ${\mathcal{S}}_{m_{x},m_{z}}\triangleq[m_{x}d-{\sqrt{A}}/{2},m_{x}d+{\sqrt{A}}/{2}]\times[m_{z}d-{\sqrt{A}}/{2},m_{z}d+{\sqrt{A}}/{2}]$ denotes the surface region of the $(m_{x},m_{z})$th array element. The term $\frac{{\mathbf{e}}_{y}^{\mathsf{T}}({\mathbf{r}}_{k}-{\mathbf{s}})}{\lVert{\mathbf{r}}_{k}-{\mathbf{s}}\rVert}$ models the influence of the projected aperture of each array element, which is reflected by the projection of the UPA normal vector ${\mathbf{e}}_{y}=[0,1,0]^{\mathsf{T}}$ onto the wave propagation direction at each local point $\mathbf{s}$ [4]. Furthermore, ${\mathsf{g}}(\mathbf{r},{\mathbf{s}})$ models the effect of free-space EM propagation, which is given by [18]
@@
with $k_{0}=\frac{2\pi}{\lambda}$ being the wavenumber. Given the small antenna size compared to the propagation distance between the user and the antenna elements, i.e., $\sqrt{A}\ll\lVert{\mathbf{r}}_{k}-{\mathbf{s}}_{m_{x},m_{z}}\rVert$, the variation in the channel response across an individual antenna element can be considered negligible. It follows from the fact of $\int_{{\mathcal{S}}_{m_{x},m_{z}}}{\rm{d}}{\mathbf{s}}=A$ that
$g_{m_{x},m_{z},k}$ $\approx{\mathsf{g}}(\mathbf{r}_{k},{\mathbf{s}}_{m_{x},m_{z}})\frac{\lvert{\mathbf{e}}_{y}^{\mathsf{T}}({\mathbf{s}}_{m_{x},m_{z}}-{\mathbf{r}}_{k})\rvert}{\lVert{\mathbf{r}}_{k}-{\mathbf{s}}_{m_{x},m_{z}}\rVert}\int_{{\mathcal{S}}_{m_{x},m_{z}}}{\rm{d}}{\mathbf{s}}$
$=A{\mathsf{g}}(\mathbf{r}_{k},{\mathbf{s}}_{m_{x},m_{z}})\frac{\lvert{\mathbf{e}}_{y}^{\mathsf{T}}({\mathbf{s}}_{m_{x},m_{z}}-{\mathbf{r}}_{k})\rvert}{r_{m_{x},m_{z},k}}.$ (9)
@@ II-B2 USW Model
Since the uniform-power distance is typically much smaller than the Rayleigh distance [4], the USW is effective for the majority of the area within the near-field region. Furthermore, as discussed in [4], when the communication distance exceeds the Fresnel distance $0.5\sqrt{\frac{D^{3}}{\lambda}}$ [19], (6b) can be approximated as follows:
@@
$\begin{split}\mathbf{h}_{k}\approx\sqrt{{A\Psi_{k}}/({4\pi r_{k}^{2}})}\left[\mathrm{e}^{-\mathrm{j}\frac{2\pi}{\lambda}\tilde{r}_{m_{x},m_{z},k}}\right]_{\forall m_{x},m_{z}}.\end{split}$ (13)
@@ II-B3 UPW Model
In the following sections, we aim to analyze (4) and (5) under the three models discussed above.
@@ III-A Secrecy Capacity Characterization
According to the definition in (4), the optimal beamformer $\mathbf{w}^{\star}$ that achieves the secrecy capacity satisfies
@@
${\mathbf{v}^{\star}}=\operatorname*{argmax}\nolimits_{\lVert{\mathbf{v}}\rVert^{2}=1}\left(\frac{1+{\overline{\gamma}_{\mathsf{b}}}\lvert{\mathbf{h}}_{\mathsf{b}}^{\mathsf{H}}{\mathbf{v}}\rvert^{2}}{1+{\overline{\gamma}_{\mathsf{e}}}\lvert{\mathbf{h}}_{\mathsf{e}}^{\mathsf{H}}{\mathbf{v}}\rvert^{2}}=\frac{{\mathbf{v}}^{\mathsf{H}}{\mathbf{Q}}_{\mathsf{b}}{\mathbf{v}}}{{\mathbf{v}}^{\mathsf{H}}{\mathbf{Q}}_{\mathsf{e}}{\mathbf{v}}}\right)$ (16)
@@
$\frac{{\rm{d}}g_{\mathsf{n}}(\rho)}{{\rm{d}}\rho}=-{\overline{\gamma}_{\mathsf{b}}}{\overline{\gamma}_{\mathsf{e}}}G_{\mathsf{b}}G_{\mathsf{e}}\left(1+\frac{\alpha}{\sqrt{\alpha^{2}+4\beta}}\right)-\frac{1+{\overline{\gamma}_{\mathsf{e}}}G_{\mathsf{e}}}{\sqrt{\alpha^{2}+4\beta}}.$

### v2

@@ I Introduction
Recent research trends have highlighted the advantages of multiple-antenna technologies in improving secrecy capacity through beamforming design [2, 4]. Intuitively, using a larger-aperture antenna array can provide a higher beamforming gain and, consequently, a higher secure transmission rate. With advancements in antenna manufacturing and to meet the stringent security demands of future networks, there is a growing trend for base stations (BSs) to employ extremely large-aperture antenna arrays and operate at significantly higher frequencies [5, 6, 7, 8]. However, this shift represents not just a quantitative increase in antenna size and carrier frequency but also a qualitative paradigm shift from conventional far-field communications (FFC) to near-field communications (NFC). Specifically, based on the criterion for differentiating between near-field and far-field regions—the Rayleigh distance $\frac{2D^{2}}{\lambda}$, where $D$ represents the antenna aperture and $\lambda$ is the wavelength [9]—the near-field region is expected to expand substantially. For instance, an array with an aperture of $D=0.5$ m operating at a frequency of 60 GHz would encompass a near-field region extending up to $100$ meters.
@@ I-A Prior Works
The secrecy performance for FFC, where the UPW model is applied, has been widely studied; see [2, 4, 11, 12, 13], and related references. However, research on near-field physical layer security is still in its early stages. The authors in [14] utilized the USW model to study a near-field secure directional modulation system and propose a secure precoding algorithm. The work in [15] leveraged the NUSW model to develop a two-stage algorithm that maximizes the near-field secrecy capacity via jointly optimizing unit-modulus phase shifters and baseband digital beamformers. The authors of [16] introduced analog beamfocusing techniques using true-time delayers and phase shifters to effectively mitigate beamsplit and exploit near-field advantages, significantly improving physical layer security in near-field wideband communications. Meanwhile, the study in [17] proposed an efficient low-complexity strategy for designing the beamforming with artificial noise to ensure near-field secure transmission with multiple legitimate users.
@@
I-B Motivations & Contributions
In light of the above discussion, the UPW model, commonly utilized in far-field scenarios, is inapplicable for NFC. Therefore, it becomes imperative to reevaluate the performance for physical layer security from a near-field perspective. By examining the propagation differences between near-field and far-field channel models and their impact on secure transmission, we aim to provide a straightforward comparison of secrecy performance between NFC and FFC, which motivates this work.
This article analyzes the performance of physical layer security across commonly used far-field and near-field channel models, namely the UPW, USW, and NUSW models. We derive closed-form expressions for both the secrecy capacity and the minimum required power. Besides, asymptotic analysis is performed to provide deeper insights. Our main contributions are summarized as follows.
  • We investigate secure transmission scenarios where a large‐aperture array communicates with a legitimate user while being eavesdropped by another user. We provide detailed analysis of various channel models applicable across different ranges in both near-field and far-field regions, including the UPW, USW and NUSW models. In this context, we introduce two key metrics to measure secrecy performance: the secrecy capacity under a power constraint and the minimum power requirement needed to achieve a target secrecy rate.
@@ II System Model
We consider a multiple-input single-output (MISO) wiretap channel, which comprises a base station (BS), a legitimate user (Bob), and an eavesdropper (Eve), as depicted in Figure 2. The BS is equipped with a large-aperture UPA containing $M\gg 1$ antenna elements, while both users are single-antenna devices  [^1] The two-user wiretap channel considered in this work is a basic model for secret communications, which has been extensively studied in prior research, e.g., [20, 21, 22, 14, 15, 16, 18, 19]. This model provides valuable insights into the fundamental limits of physical layer security and has facilitated significant theoretical advancements in secure communications, as demonstrated in [23, 21, 22, 20, 24, 3]. Since this research represents an initial exploration into the fundamental performance limits of near-field physical layer security, we focus on this basic model to derive deeper insights into the system. Future studies will extend this analysis to more complex scenarios with multiple users.. The UPA is placed on the $x$-$z$ plane and centered at the origin. Here, $M=M_{x}M_{z}$, where $M_{x}$ and $M_{z}$ denote the number of array elements along the $x$- and $z$-axes, respectively. For brevity, we assume that $M_{x}$ and $M_{z}$ are odd numbers with $M_{x}=2\tilde{M}_{x}+1$ and $M_{z}=2\tilde{M}_{z}+1$. The physical dimensions of each BS array element along the $x$- and $z$-axes are denoted by $\sqrt{A}$, and the inter-element distance is $d$, where $d\geq\sqrt{A}$.
@@ II-A1 Secrecy Channel Capacity
${\mathsf{C}}$ $=\max_{\lVert{\mathbf{w}}\rVert^{2}\leq P}{\mathsf{C}}_{\mathsf{w}}$
$=\max_{\lVert{\mathbf{w}}\rVert^{2}\leq P}\left(\max\left\{\log_{2}\left(\frac{1+\gamma_{\mathsf{b}}}{1+\gamma_{\mathsf{e}}}\right),0\right\}\right).$ (4)
@@ II-B Channel Model
All channel gains are fixed throughout the entire transmission period, and each node $k\in\{{\mathsf{b}},{\mathsf{e}}\}$ is assumed to possess complete channel state information (CSI) about its effective channel. It is further assumed that Eve is a registered user, and hence, during the channel training phase, the BS acquires the CSI of both Eve and Bob. Under these circumstances, Eve is expected to receive the common messages broadcasted in the network but should remain uninformed about the confidential messages intended for Bob  [^2] This assumption is widely used in the physical layer security literature because many key secrecy performance metrics, such as secrecy channel capacity [23], and foundational results, including secrecy channel coding theorems [23, 21, 22, 3], are derived based on it. Consequently, numerous studies on near-field physical layer security have also adopted this classical model; see [15] and related references. For scenarios where Eve operates as a malicious node with unknown CSI and multiple antennas, techniques like artificial noise can be employed to enhance security; detailed discussions can be found in [25, 26, 27]. However, addressing such cases falls outside the scope of this work and will be explored in future research..
Since near-field channels are sparsely-scattered and dominated by line-of-sight (LoS) propagation [5], we consider pure LoS propagation scenarios for a theoretical exploration of fundamental secrecy performance limits and variations as the propagation transitions from far-field to near-field. Following our description concerning Figure 2, the central location of the $(m_{x},m_{z})$th BS element is denoted by ${\mathbf{s}}_{m_{x},m_{z}}=[m_{x}d,0,m_{z}d]^{\mathsf{T}}$, where $m_{x}\in{\mathcal{M}}_{x}\triangleq\{0,\pm 1,\ldots,\pm\tilde{M}_{x}\}$ and $m_{z}\in{\mathcal{M}}_{z}\triangleq\{0,\pm 1,\ldots,\pm\tilde{M}_{z}\}$. Let $r_{k}$ denote the distance from the center of the antenna array to each node $k\in\{{\mathsf{b}},{\mathsf{e}}\}$, and $\theta_{k}\in(0,\pi)$ and $\phi_{k}\in(0,\pi)$ denote the azimuth and elevation angles, respectively. Thus, the location of node $k$ can be expressed as ${\mathbf{r}}_{k}=[r_{k}\Phi_{k},r_{k}\Psi_{k},r_{k}\Omega_{k}]^{\mathsf{T}}$, where $\Phi_{k}\triangleq\sin{\phi_{k}}\cos{\theta_{k}}$, $\Psi_{k}\triangleq\sin\phi_{k}\sin\theta_{k}$, and $\Omega_{k}\triangleq\cos{\phi_{k}}$. Moreover, we assume that both Bob and Eve are equipped with a single hypothetical isotropic antenna to receive incoming signals.
@@ II-B1 NUSW Model
$r_{m_{x},m_{z},k}=\lVert{\mathbf{r}}_{k}-{\mathbf{s}}_{m_{x},m_{z}}\rVert$
$=r_{k}\sqrt{1-2m_{x}\varepsilon_{k}\Phi_{k}-2m_{z}\varepsilon_{k}\Omega_{k}+(m_{x}^{2}+m_{z}^{2})\varepsilon_{k}^{2}},$ (6)
@@
$g_{m_{x},m_{z},k}=\int_{{\mathcal{S}}_{m_{x},m_{z}}}{\frac{\lvert{\mathbf{e}}_{y}^{\mathsf{T}}({\mathbf{s}}-{\mathbf{r}}_{k})\rvert}{\lVert{\mathbf{r}}_{k}-{\mathbf{s}}\rVert}}{\mathsf{g}}(\mathbf{r}_{k},{\mathbf{s}}){\rm{d}}{\mathbf{s}}.$ (7)
where ${\mathcal{S}}_{m_{x},m_{z}}\triangleq[m_{x}d-{\sqrt{A}}/{2},m_{x}d+{\sqrt{A}}/{2}]\times[m_{z}d-{\sqrt{A}}/{2},m_{z}d+{\sqrt{A}}/{2}]$ denotes the surface region of the $(m_{x},m_{z})$th array element. The term $\frac{{\mathbf{e}}_{y}^{\mathsf{T}}({\mathbf{r}}_{k}-{\mathbf{s}})}{\lVert{\mathbf{r}}_{k}-{\mathbf{s}}\rVert}$ models the influence of the projected aperture of each array element, which is reflected by the projection of the UPA normal vector ${\mathbf{e}}_{y}=[0,1,0]^{\mathsf{T}}$ onto the wave propagation direction at each local point $\mathbf{s}$ [5]. It is applied to individual array elements that explicitly accounts for their size. Furthermore, ${\mathsf{g}}(\mathbf{r},{\mathbf{s}})$ models the effect of free-space EM propagation, which is given by [28]
@@
with $k_{0}=\frac{2\pi}{\lambda}$ being the wavenumber.
Given the small antenna size compared to the propagation distance between the user and the antenna elements, i.e., $\sqrt{A}\ll\lVert{\mathbf{r}}_{k}-{\mathbf{s}}_{m_{x},m_{z}}\rVert$, the variation in the channel response across an individual antenna element can be considered negligible. It follows from the fact of $\int_{{\mathcal{S}}_{m_{x},m_{z}}}{\rm{d}}{\mathbf{s}}=A$ that
$g_{m_{x},m_{z},k}$ $\approx\frac{\lvert{\mathbf{e}}_{y}^{\mathsf{T}}({\mathbf{s}}_{m_{x},m_{z}}-{\mathbf{r}}_{k})\rvert}{\lVert{\mathbf{r}}_{k}-{\mathbf{s}}_{m_{x},m_{z}}\rVert}{\mathsf{g}}(\mathbf{r}_{k},{\mathbf{s}}_{m_{x},m_{z}})\int_{{\mathcal{S}}_{m_{x},m_{z}}}{\rm{d}}{\mathbf{s}}$
$=A\frac{\lvert{\mathbf{e}}_{y}^{\mathsf{T}}({\mathbf{s}}_{m_{x},m_{z}}-{\mathbf{r}}_{k})\rvert}{r_{m_{x},m_{z},k}}{\mathsf{g}}(\mathbf{r}_{k},{\mathbf{s}}_{m_{x},m_{z}}).$ (9)
@@ II-B2 USW Model
Since the uniform-power distance is typically much smaller than the Rayleigh distance [5], the USW is effective for the majority of the area within the near-field region. Furthermore, as discussed in [5], when the communication distance exceeds the Fresnel distance $0.5\sqrt{\frac{D^{3}}{\lambda}}$ [29], (6) can be approximated as follows:
@@
$\begin{split}\mathbf{h}_{k}\approx\sqrt{\frac{A\Psi_{k}}{4\pi r_{k}^{2}}}\left[\mathrm{e}^{-\mathrm{j}\frac{2\pi}{\lambda}\tilde{r}_{m_{x},m_{z},k}}\right]_{\forall m_{x},m_{z}}.\end{split}$ (13)
@@ II-B3 UPW Model
In the following sections, we aim to analyze (II-A1) and (5) under the three models discussed above.
@@ III-A Secrecy Capacity Characterization
According to the definition in (II-A1), the optimal beamformer $\mathbf{w}^{\star}$ that achieves the secrecy capacity satisfies
@@
${\mathbf{v}^{\star}}=\operatorname*{argmax}\nolimits_{\lVert{\mathbf{v}}\rVert^{2}=1}\left(\frac{1+{\overline{\gamma}_{\mathsf{b}}}\lvert{\mathbf{h}}_{\mathsf{b}}^{\mathsf{H}}{\mathbf{v}}\rvert^{2}}{1+{\overline{\gamma}_{\mathsf{e}}}\lvert{\mathbf{h}}_{\mathsf{e}}^{\mathsf{H}}{\mathbf{v}}\rvert^{2}}=\frac{{\mathbf{v}}^{\mathsf{H}}{\mathbf{Q}}_{\mathsf{b}}{\mathbf{v}}}{{\mathbf{v}}^{\mathsf{H}}{\mathbf{Q}}_{\mathsf{e}}{\mathbf{v}}}\right),$ (16)
@@
$\frac{{\rm{d}}g_{\mathsf{n}}(\rho)}{{\rm{d}}\rho}=-{\overline{\gamma}_{\mathsf{b}}}{\overline{\gamma}_{\mathsf{e}}}G_{\mathsf{b}}G_{\mathsf{e}}\left(1+\frac{\alpha}{\sqrt{\alpha^{2}+4\beta}}\right)-\frac{1+{\overline{\gamma}_{\mathsf{e}}}G_{\mathsf{e}}}{\sqrt{\alpha^{2}+4\beta}}.$ (22)